OpenStudy (anonymous):

What is the radius of the circle whose equation is x2 + y2 = 8?

OpenStudy (anonymous):

\[ x^2+y^2=r^2 \] r = radius (take the square root)

OpenStudy (anonymous):

8 = 2 * 2 * 2 Does this answer your question? :-)